OpenStudy (anonymous):

solve the following simultaneous equations using graphical method: (7) x+4y=12 4x+y=18 Qno.(4) 4x+y=2 4x+y=-3

OpenStudy (anonymous):

@dan815 @alekos @TheSmartOne @freckles @uri @OwlGirl212 @iambatman

OpenStudy (anonymous):

are you solving the system of equations or solving for x?

OpenStudy (anonymous):

@Nurali @nincompoop @nickmeneses1997 @Emmanuel11 @Scotty4life

OpenStudy (anonymous):

for find the value of x and y?

OpenStudy (anonymous):

it looks like you have two questions there, is that right?

OpenStudy (anonymous):

yes 4 and 7 first i want to solve first one...!

OpenStudy (anonymous):

ok so to solve this graphically you need to do it by hand or use a graphing calculator, graph both lines, at the point where they intersect is the solution it will be at some (x,y) coordinate that they intersect

OpenStudy (anonymous):

are you allowed to use a graphing calculator?

OpenStudy (anonymous):

no .. i want to solve without calculator

OpenStudy (anonymous):

ok so can you graph those lines?

OpenStudy (anonymous):

by hand?

OpenStudy (anonymous):

yes by hand?

OpenStudy (anonymous):

yes graph the lines by hand, observe what (x,y) coordinate they intersect at, that is the solution

OpenStudy (anonymous):

first we find the value of x and y,,,,,how?

OpenStudy (wolf1728):

Here's a graph if you need one:

OpenStudy (anonymous):

it may be easiest if you convert the equations to slope intercept form, y = mx+b before graphing

OpenStudy (anonymous):

you actually dont need to though

OpenStudy (anonymous):

give me an x value and a y value that will make this equation true x + 4y = 12

OpenStudy (anonymous):

ill give you one, how about x = 8 and y = 1, 8 + 4(1) = 12

OpenStudy (anonymous):

can you give me another one?

OpenStudy (anonymous):

i want to retell the first one in this way.... if x=0 and y=12/4=3

OpenStudy (anonymous):

not sure what you mean by that

OpenStudy (anonymous):

@wolf1728 i make this question by graph calculator as give below oooops.....my copy pasting is not working

OpenStudy (anonymous):

x +4y =12 we find the value of y , whats the answer of y?

OpenStudy (anonymous):

oh so you want to solve for y?

OpenStudy (anonymous):

ok so you want to isolate y on one side of the equation

OpenStudy (anonymous):

yes for x and y..

OpenStudy (anonymous):

@wolf my computer is not working as pasting???

OpenStudy (wolf1728):

ziqball103 You have to use the attach file button

OpenStudy (anonymous):

doing copy but not pasting? i want to show you graph calculator how i paste link ???

OpenStudy (anonymous):

@ziqbal103 if you want to solve for x and y for that system graphically, you must graph them, thats what graphically means

OpenStudy (anonymous):

i thing my windows dissapear that's why i cann't paste my link @ wolf

OpenStudy (wolf1728):

You are trying to paste a link? CAn you just graph the equation on your own computer?

OpenStudy (anonymous):

@billj5 find the value of x and y

OpenStudy (anonymous):

@ziqbal103 i dont know what to tell you, you arent listening to what im saying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ziqbal103 if you come back, msg me and we can work through this problem